OpenStudy (anonymous):

how do you find the antiderivative of 1/7x

OpenStudy (anonymous):

it is given by formula is \[\Large \int\limits_{}^{} \frac{1}{x}=\ln(x)+C\] so in the given question \[\Large \int\limits_{}^{}\frac{1}{7x}=\frac{1}{7}\int\limits_{}^{}\frac{1}{x}=\frac{1}{7}\ln(x)+C\]

OpenStudy (anonymous):

i missed dx !!!!

OpenStudy (anonymous):

@fheard did you get it?

OpenStudy (anonymous):

@fheard waiting for your response !

OpenStudy (anonymous):

sorry started another problem yes I got it thanks